OpenStudy (anonymous):

Find the circumference. diameter=25.8

OpenStudy (anonymous):

r u that dumb bruv lol

OpenStudy (anonymous):

\[C = \pi d\]

OpenStudy (gorv):

circumference=2*pi*r diameter= 2*r r=25.8/2=12.9

OpenStudy (anonymous):

C=Circumference d=diameter

OpenStudy (anonymous):

k